A 1.34 g sample of caffeine (C8H10N4O2) burns in a constant-volume calorimeter that has a heat capacity of 7.71 kJ/K. The temperature increases from 297.65 K to 302.07 K. Determine the heat (qv) associated with this reaction. 1476 kJ/mol

Now use the data above to find ΔE for the combustion of one mole of caffeine.
Answer #1

qv = mSdT

= 7.71 KJ/K * (302.07-297.65)

= 34.078 KJ

Now mole of caffeine present = 1.34/ 194 = 6.9*10^-3 moles

so heat per mole = 34.078/ 6.9*10^-3 = 4933.68 KJ/mole